A computer store manager buys several computers of the same model for $10,800. The store can regain this investment by selling all but 5 of the computers at a profit of $360 per computer. To do this, how many computers must be sold, and at what price?
______computers at $_____ per computer

Answer by stanbon(75887)

A computer store manager buys several computers of the same model for $10,800. The store can regain this investment by selling all but 5 of the computers at a profit of $360 per computer. To do this, how many computers must be sold, and at what price?
# of computers bought:: "x"
cost of each computer:: 10800/x
------
Equation: (x-5)[(10800/x)+360] = 10,800
------
[10800+ 360x-5(10800/x)-1800] = 10,800
------
360x - 54000/x -1800 = 0
-54000 - 1800x + 360x^2 = 0
x^2 - 5x - 150 = 0
(x-15)(x+10) = 0
---
Positive solution:
x = 15
-----
Ans:: x-5 = 15-5 = 10 must be sold
----
Selling price when x = 15
[(10800/x)+360] = 720+360 = $1080
-------------
Cheers,
Stan H.
-------------




______computers at $_____ per computer